# Infer-and-widen versus split-and-condition: two tales of selective inference

### v1

@@ A.2 Proof of Proposition 1
Appendix B Vignette #2
B.1 Proof of Proposition 2
@@
Let $Y\sim{\bf{\rm N}}\,_{n}(\mu,\Sigma)$ an $n$-dimensional random vector, and $\beta\in\mathbb{R}^{n}$ a vector of interest. For any fixed matrix $D\in\mathbb{R}^{n\times d}$ and vector $b\in\mathbb{R}^{d}$, define
@@
$\beta^{\top}Y|\left(DY\leq b,W\right)\sim{\bf{\rm N}}\,_{[v_{\min},v_{\max}]}(\beta^{\top}\mu,\beta^{\top}\Sigma\beta),$ (B.1)
@@
Returning to our setting under the model in 3.1 with a fixed and column-normalized design matrix $X$, for brevity let $Z:=X^{\top}Y$ and $Z^{\mathrm{train}}:=Z+\zeta$ where $\zeta_{1},\ldots,\zeta_{p}\stackrel{{\scriptstyle{i.i.d.}}}{{\sim}}{\bf{\rm Laplace}}\,(c)$. Recall that our target of inference is $X_{\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(Y)}^{\top}\mu$, where $\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(Y)$ is the selection rule in 4.2. This selection rule is implicitly a function of the Laplacian noise $\zeta$, and so to emphasize this point we will write $\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(Y;\zeta)$ in this section.
To construct a confidence interval, as in 2.6 it will suffice to derive the distribution of $X_{\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(Y;\zeta)}^{\top}Y\mid C(Y,\zeta)$ where $\sigma\left(C(Y,\zeta)\right)\supseteq\sigma\left(\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(Y;\zeta)\right)$, e.g. we will condition on the selection event along with an additional event. By the law of total probability, conditional coverage implies marginal coverage.
@@
$\left\{(Y,\zeta):\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(Y;\zeta)=\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(y;\zeta^{obs})\right\}=\left\{(Y,\zeta):\delta A_{\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(y;\zeta^{obs})}X^{\top}Y\leq-\delta A_{\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(y;\zeta^{obs})}\zeta\right\},$ (B.2)
@@
$\delta:={\rm sign}\left(X_{\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(y;\zeta^{obs})}^{\top}Y+\zeta_{\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(y;\zeta^{obs})}\right)={\rm sign}\left(Z^{\mathrm{train}}_{\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(y;\zeta^{obs})}\right).$ (B.3)
@@
To show that there exists a matrix $A$ for which (B.2) holds, note that by definition, $\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(Y;\zeta)=\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(y;\zeta^{obs})$ if and only if $|Z^{\mathrm{train}}_{\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(y;\zeta^{obs})}|\geq|Z^{\mathrm{train}}_{j}|$ for all $j\in[p]$. In turn, expanding the absolute value and definition of $Z^{\mathrm{train}}$, (B.2) holds if and only if both (I) and (II) hold:
@@
$A:=\begin{bmatrix}0_{1\times n_{1}}&-1&0_{1\times n_{2}}\\ I_{n_{1}}&-1_{n_{1}}&0_{n_{1}\times n_{2}}\\ 0_{n_{2}\times n_{1}}&-1_{n_{2}}&I_{n_{2}}\\ -I_{n_{1}}&-1_{n_{1}}&0_{n_{1}\times n_{2}}\\ 0_{n_{2}\times n_{1}}&-1_{n_{2}}&-I_{n_{2}}\end{bmatrix},$ (B.4)
and we denote $n_{1}:=\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(y;\zeta^{obs})-1$ and $n_{2}:=p-\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(y;\zeta^{obs})$ for brevity. If instead $Z^{\mathrm{train}}_{\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(y;\zeta^{obs})}\leq 0$, then we are in case (II), where $\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(Y;\zeta)=\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(y;\zeta^{obs})$ if and only if $Z$ satisfies $-AZ\leq A\zeta$. Thus, in either case, $\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(Y;\zeta)=\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(y;\zeta^{obs})$ if and only if $\delta AZ\leq-\delta A\zeta$, and so, by definition of $Z$, if and only if $\delta AX^{\top}Y\leq-\delta A\zeta$. This establishes that (B.2) holds for $A$ defined in (B.4).
Thus, by B.2, we see that
@@
$X_{\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(Y;\zeta)}^{\top}Y|\left(\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(Y;\zeta)=\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(y;\zeta^{obs}),\zeta=\zeta^{obs},\delta=\delta^{obs},W\right)$ (B.5)
@@
=$\stackrel{{\scriptstyle\mathrm{d}}}{{=}}X_{\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(y;\zeta^{obs})}^{\top}Y|\left(\delta AX^{\top}Y\leq-\delta A\zeta,\zeta=\zeta^{obs},\delta=\delta^{obs},W\right)$ (B.6)
@@
=$\stackrel{{\scriptstyle\mathrm{d}}}{{=}}X_{\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(y;\zeta^{obs})}^{\top}Y|\left(\delta^{obs}AX^{\top}Y\leq-\delta^{obs}A\zeta,\zeta=\zeta^{obs},W\right)$ (B.7)
@@
=$\stackrel{{\scriptstyle\mathrm{d}}}{{=}}X_{\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(y;\zeta^{obs})}^{\top}Y|\left(\delta^{obs}AX^{\top}Y\leq-\delta^{obs}A\zeta^{obs},W\right).$ (B.8)
B.6 follows from the equivalence in B.2.
To establish B.7, note that the event $\{\delta^{obs}AX^{\top}Y\leq-\delta^{obs}A\zeta\}$ is composed of a system of $12p+$ 1 inequalities, the first of which states that $-\delta^{obs}X_{\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(y;\zeta^{obs})}^{\top}Y\leq\delta^{obs}\zeta_{\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(y;\zeta^{obs})}$, or equivalently, that $0\leq\delta^{obs}\left(X_{\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(y;\zeta^{obs})}^{\top}Y+\zeta_{\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(y;\zeta^{obs})}\right)$. Recalling the definition of $\delta$ in (B.3), it follows that the event $\{\delta^{obs}AX^{\top}Y\leq-\delta^{obs}A\zeta\}$ implies that $\delta=\delta^{obs}$. This yields (B.7).
@@
$X_{\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(y;\zeta^{obs})}^{\top}Y|\left(\delta^{obs}AX^{\top}Y\leq-\delta^{obs}A\zeta^{obs},W\right)\sim{\bf{\rm N}}\,_{[v_{\min},v_{\max}]}(X_{\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(y;\zeta^{obs})}^{\top}\mu,\sigma^{2}),$ (B.9)
for $v_{\min}$ and $v_{\max}$ given in the statement of Proposition 2, which are functions of $\zeta^{obs}$ and $\delta^{obs}$ through $D$ and $b$.
Applying B.5-B.8 to (B.9), we have that
@@
A confidence interval for $X_{\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(Y;\zeta)}^{\top}\mu~{}|\left(\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(Y;\zeta),\zeta,\delta,W\right)$ is thus obtained using the CDF of the above truncated normal distribution. This in turn provides valid unconditional coverage for $X_{\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(Y;\zeta)}^{\top}\mu$ by the law of total probability. ∎
B.2 Sample splitting versus data thinning
Not all split-and-condition methods are alike. For instance, at first glance, it may appear that sample splitting and data thinning [Rasines and Young, 2023, Neufeld et al., 2024, Dharamshi et al., 2024] are comparable: both approaches yield independent training and test sets, and the widths of their confidence intervals depend on the fraction of the Fisher information in the test set. However, the two methods can fundamentally differ wildly. This explains why our results for Vignette #2 in Section 4 comparing algorithmic stability to split-and-condition approaches other than sample splitting, with identical selection events, differ from the results seen in Figure 2 of Zrnic and Jordan [2023].
Consider the setup in Section 4 with $n=p$, and let $X=I_{p}$ and $Y\sim{\bf{\rm N}}\,_{p}(\mu,I_{p})$. Suppose we conduct sample splitting and put $\alpha\in[0,1]$ fraction of the observations into the training sets $X_{\alpha}^{\mathrm{train}},Y_{\alpha}^{\mathrm{train}}$, and the rest into the test sets $X_{\alpha}^{\mathrm{test}},Y_{\alpha}^{\mathrm{test}}$. Let $j^{*}:=\mathop{\mathrm{arg\,max}}_{j}\left|(X_{\alpha}^{\mathrm{train}})^{\top}_{j}Y_{\alpha}^{\mathrm{train}}\right|$. Then, $\Pr\left(j^{*}=\mathop{\mathrm{arg\,max}}_{j}\left|X^{\top}_{j}Y\right|\right)=\alpha$ since with probability $\alpha$ the training set contains the row of $X$ corresponding to the maximal magnitude element of $Y$. Note that this probability is independent of $\mu$. Furthermore, the $j^{*}$th column in $X_{\alpha}^{\mathrm{test}}$ is necessarily a column of zeros, since $X$ is the identity and the $j^{*}$th column of $X_{\alpha}^{\mathrm{train}}$ had a one in it. Thus, $(X_{\alpha}^{\mathrm{test}})^{\top}_{j^{*}}Y_{\alpha}^{\mathrm{test}}$ is a degenerate distribution with mean 00 and variance 00.
In contrast, data thinning constructs independent folds $Y^{\mathrm{train}}=Y+c\zeta$ and $Y^{\mathrm{test}}=Y-\tfrac{1}{c}\zeta$, $\zeta_{i}\stackrel{{\scriptstyle{i.i.d.}}}{{\sim}}{\bf{\rm N}}\,(0,1)$. Let $j^{*}:=\mathop{\mathrm{arg\,max}}_{j}\left|X^{\top}_{j}Y^{\mathrm{train}}\right|$. Then, $\Pr\left(j^{*}=\mathop{\mathrm{arg\,max}}_{j}\left|X^{\top}_{j}Y\right|\right)$ occurs with probability dependent on $p$, $\mu$, and $c$. Furthermore, $X_{j^{*}}^{\top}Y^{\mathrm{test}}\sim{\bf{\rm N}}\,(\mu_{j^{*}},1+\tfrac{1}{c^{2}})$, leading to a non-degenerate distribution. We see that both selection and inference differ substantially from sample splitting in this case.
B.3 Additional simulation results
The simulation results in Section 4 reflect a limited set of parameter choices. Figure 4 extends Figure 2(a) to a wider range of levels of signal. Figure 5 extends Figures 2(b, c) to a wider range of levels of signal.
@@ B.3 Additional simulation results
Figure 4: Vignette #2, under the model (3.1) with $n=100$, $\sigma=1$, $\mu=X^{\top}\phi$, and selection rule in 4.2, where $X$ is multivariate normal with column correlation $0.5$, and $\phi$ is sparse with exponentially distributed non-zero elements. We vary the mean of $\phi$ across $0,5,50/7$ in panels (a), (b), and (c), respectively. Results are averaged over $250$ simulated datasets. We display the ratio of widths of infer-and-widen 4.3 and split-and-condition 4.5 intervals with $95\%$ coverage across the number of features $p$ and Laplacian noise variance $2c^{2}$. A width ratio that exceeds one implies that the infer-and-widen interval is wider. Black regions represent pairs of $n$ and $c$ for which the infer-and-widen approach requires such a small Gaussian quantile that R returns an interval of infinite width. It seems that the differences in widths is most significant under the null in panel (a), but results are qualitatively similar across levels of signal.
@@
Figure 5: Vignette #2, under the model (3.1) with $n=100$, $p=100$, $\sigma=1$, $\mu=X^{\top}\phi$, and selection rule in 4.2, where $X$ is multivariate normal with column correlation $0.5$, and $\phi$ is sparse with exponentially distributed non-zero elements. We vary the mean of $\phi$ across the three columns, and the Laplacian noise variance across the two rows. We display the width of the “oracle” infer-and-widen interval and the split-and-condition interval 4.5 as a function of coverage. The 95% classical interval, and the “attainable” 95% infer-and-widen interval 4.3 based on algorithmic stability, are also displayed. Results are qualitatively similar regardless of choice of parameters.
Appendix C Vignette #3
C.1 Proof of LABEL:{thm:vignette-3-gaussian-DT}

### v2

@@ B.2 Proof of Proposition 6
Appendix C Vignette #2
C.1 Proof of Proposition 4
@@
Let $Y\sim{\bf{\rm N}}\,_{n}(\mu,\Sigma)$ be an $n$-dimensional random vector, and $\beta\in\mathbb{R}^{n}$ a vector of interest. For any fixed matrix $D\in\mathbb{R}^{n\times d}$ and vector $b\in\mathbb{R}^{d}$, define
@@
$\beta^{\top}Y|\left(DY\leq b,W\right)\sim{\bf{\rm N}}\,_{[v_{\min},v_{\max}]}(\beta^{\top}\mu,\beta^{\top}\Sigma\beta),$ (C.1)
@@
Returning to our setting in Proposition 4, for brevity let $Z:=X^{\top}Y$ and $Z^{\mathrm{train}}:=Z+\zeta$ where $\zeta_{1},\ldots,\zeta_{p}\stackrel{{\scriptstyle{i.i.d.}}}{{\sim}}{\bf{\rm Laplace}}\,(c)$. Recall that our target of inference is $X_{\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(Y)}^{\top}\mu$, where $\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(Y)$ is the selection rule in (4.2). This selection rule is implicitly a function of the Laplacian noise $\zeta$, and so to emphasize this point we will write $\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(Y;\zeta)$ in this section.
To construct a confidence interval with conditional coverage as in (1.3), it will suffice to derive the distribution of $X_{\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(Y;\zeta)}^{\top}Y\mid C(Y,\zeta)$ for an event $\mathcal{C}(Y)$ satisfying $\sigma\left(C(Y,\zeta)\right)\supseteq\sigma\left(\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(Y;\zeta)\right)$, where $\sigma(\cdot)$ denotes the sigma algebra generated by a random variable, e.g., we will condition on the selection event along with an additional event.
@@
$\left\{(Y,\zeta):\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(Y;\zeta)=\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(y;\zeta^{obs})\right\}=\left\{(Y,\zeta):\delta A_{\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(y;\zeta^{obs})}X^{\top}Y\leq-\delta A_{\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(y;\zeta^{obs})}\zeta\right\},$ (C.2)
@@
$\delta:={\rm sign}\left(X_{\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(y;\zeta^{obs})}^{\top}Y+\zeta_{\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(y;\zeta^{obs})}\right)={\rm sign}\left(Z^{\mathrm{train}}_{\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(y;\zeta^{obs})}\right).$ (C.3)
@@
To show that there exists a matrix $A$ for which (C.2) holds, note that by definition $\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(Y;\zeta)=\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(y;\zeta^{obs})$ if and only if $|Z^{\mathrm{train}}_{\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(y;\zeta^{obs})}|\geq|Z^{\mathrm{train}}_{j}|$ for all $j\in[p]$. In turn, expanding the absolute value and definition of $Z^{\mathrm{train}}$, (C.2) holds if and only if both (I) and (II) hold:
@@
$A:=\begin{bmatrix}0_{1\times n_{1}}&-1&0_{1\times n_{2}}\\ I_{n_{1}}&-1_{n_{1}}&0_{n_{1}\times n_{2}}\\ 0_{n_{2}\times n_{1}}&-1_{n_{2}}&I_{n_{2}}\\ -I_{n_{1}}&-1_{n_{1}}&0_{n_{1}\times n_{2}}\\ 0_{n_{2}\times n_{1}}&-1_{n_{2}}&-I_{n_{2}}\end{bmatrix},$ (C.4)
and we denote $n_{1}:=\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(y;\zeta^{obs})-1$ and $n_{2}:=p-\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(y;\zeta^{obs})$ for brevity. If instead $Z^{\mathrm{train}}_{\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(y;\zeta^{obs})}\leq 0$, then we are in case (II), where $\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(Y;\zeta)=\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(y;\zeta^{obs})$ if and only if $Z$ satisfies $-AZ\leq A\zeta$. Thus, in either case, $\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(Y;\zeta)=\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(y;\zeta^{obs})$ if and only if $\delta AZ\leq-\delta A\zeta$, and so, by definition of $Z$, if and only if $\delta AX^{\top}Y\leq-\delta A\zeta$. This establishes that (C.2) holds for $A$ defined in (C.4).
Thus, by C.2, we see that
@@
$X_{\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(Y;\zeta)}^{\top}Y|\left(\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(Y;\zeta)=\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(y;\zeta^{obs}),\zeta=\zeta^{obs},\delta=\delta^{obs},W\right)$ (C.5)
@@
=$\stackrel{{\scriptstyle\mathrm{d}}}{{=}}X_{\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(y;\zeta^{obs})}^{\top}Y|\left(\delta AX^{\top}Y\leq-\delta A\zeta,\zeta=\zeta^{obs},\delta=\delta^{obs},W\right)$ (C.6)
@@
=$\stackrel{{\scriptstyle\mathrm{d}}}{{=}}X_{\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(y;\zeta^{obs})}^{\top}Y|\left(\delta^{obs}AX^{\top}Y\leq-\delta^{obs}A\zeta,\zeta=\zeta^{obs},W\right)$ (C.7)
@@
=$\stackrel{{\scriptstyle\mathrm{d}}}{{=}}X_{\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(y;\zeta^{obs})}^{\top}Y|\left(\delta^{obs}AX^{\top}Y\leq-\delta^{obs}A\zeta^{obs},W\right).$ (C.8)
C.6 follows from the equivalence in C.2.
To establish C.7, note that the event $\{\delta^{obs}AX^{\top}Y\leq-\delta^{obs}A\zeta\}$ is composed of a system of $12p+$ 1 inequalities, the first of which states that $-\delta^{obs}X_{\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(y;\zeta^{obs})}^{\top}Y\leq\delta^{obs}\zeta_{\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(y;\zeta^{obs})}$, or equivalently, that $0\leq\delta^{obs}\left(X_{\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(y;\zeta^{obs})}^{\top}Y+\zeta_{\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(y;\zeta^{obs})}\right)$. Recalling the definition of $\delta$ in (C.3), it follows that the event $\{\delta^{obs}AX^{\top}Y\leq-\delta^{obs}A\zeta\}$ implies that $\delta=\delta^{obs}$. This yields (C.7).
@@
$X_{\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(y;\zeta^{obs})}^{\top}Y|\left(\delta^{obs}AX^{\top}Y\leq-\delta^{obs}A\zeta^{obs},W\right)\sim{\bf{\rm N}}\,_{[v_{\min},v_{\max}]}(X_{\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(y;\zeta^{obs})}^{\top}\mu,\sigma^{2}),$ (C.9)
for $v_{\min}$ and $v_{\max}$ defined by
$v_{j}$ $:=-(A\zeta+AX^{\top}W)_{j}/(AX^{\top}X_{\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(Y)})_{j},$
$v_{\max}$ $:=\underset{j:\left(\delta AX^{\top}X_{\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(Y)}\right)_{j}<0}{\mathrm{max}}v_{j},$
$v_{\min}$ $:=\min_{j:\left(\delta AX^{\top}X_{\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(Y)}\right)_{j}>0}v_{j}.$
Applying C.5-C.8 to (C.9), we have that
@@
A confidence interval for $X_{\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(Y;\zeta)}^{\top}\mu~{}|\left(\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(Y;\zeta),\zeta,\delta,W\right)$ with conditional coverage guarantee (1.3), where $\mathcal{C}(Y)=\left(\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(Y),\zeta,\delta,W\right)$, is thus obtained using the CDF of the above truncated normal distribution. ∎
C.2 Additional simulation results
The simulation results in Section 4 reflect a limited set of parameter choices. Figure A1 extends Figure 2(a) to a wider range of levels of signal. Figure A2 extends Figure 2(b, c) to a wider range of levels of signal.
@@ C.2 Additional simulation results
Figure A1: Vignette #2, under the model (3.1) with $n=100$, $\sigma=1$, $\mu=X\phi$, and the selection rule in (4.2), where $X$ is multivariate normal with column correlation $0.5$, and $\phi$ is sparse with exponentially distributed non-zero elements. We vary the mean of $\phi$ across $0,5,50/7$ in panels (a), (b), and (c), respectively. Results are averaged over $250$ simulated datasets. We display the ratio of mean widths of infer-and-widen (4.3) and randomized conditional selective inference (RCSI) (4.6) intervals with $95\%$ coverage across the number of features $p$ and Laplacian noise variance $2c^{2}$. A width ratio that exceeds one implies that the average infer-and-widen interval is wider. Black regions represent pairs of $n$ and $c$ for which the infer-and-widen approach requires such a small Gaussian quantile that R returns an interval of infinite width. The differences in widths is most pronounced under the null in panel (a), but results are qualitatively similar across levels of signal.
@@
Figure A2: Vignette #2, under the model (3.1) with $n=100$, $p=100$, $\sigma=1$, $\mu=\phi$, and the selection rule in (4.2), where $X$ is multivariate normal with column correlation $0.5$, and $\phi$ is sparse with exponentially distributed non-zero elements. We vary the mean of $\phi$ across the three columns, and the Laplacian noise variance across the two rows. We display the average width of the “oracle” infer-and-widen interval and the randomized conditional selective inference (RCSI) interval (4.6) as a function of coverage. The 95% classical interval, and the “attainable” 95% infer-and-widen interval (4.3) based on algorithmic stability, are also displayed. Results are qualitatively similar regardless of choice of parameters.
Appendix D Vignette #3
D.1 Proof of Proposition 5